# Magneto-optical trap loading with an effusive oven in    a large optical access experiment

### v1

@@ VI. Acknowledgments
We are grateful for contributions to the experimental setup by A. Rybicki and J. McDowell. This work was supported by an award from the W. M. Keck Foundation. M. Gaudesius was supported in part by the Dodge Family Foundation, and L. Kraft was supported in part by the Avenir Foundation.
Appendix A1
@@
We base our model on the hyperfine transition $F=0\rightarrow F^{\prime}=1$, faithful to the blue (or red) trapping descriptions of 88Sr. Here, each of the three Zeeman transitions $m_{0}=0\rightarrow m_{1}=-1,0,+1$ between the hyperfine levels is treated as an independent 2-level system and driven by, respectively, $\sigma^{-}$, $\pi$, $\sigma^{+}$ polarized light (refer to Fig. 1(b) in Ref. [54]). This approximation is expected to hold only in the regime of low saturation, $s=\frac{I_{L}/I_{sat}}{1+\frac{4\Delta^{2}}{\Gamma^{2}}}\ll 1$, where $I_{L}$ is the laser intensity, $I_{sat}$ is the saturation intensity, $\Delta=\omega_{L}-\omega_{0}$ is the detuning of the laser frequency $\omega_{L}$ from the atomic transition $m_{0}=0\rightarrow m_{1}=0$ frequency $\omega_{0}$, and $\Gamma$ is the natural linewidth of the hyperfine transition.
We take into account arbitrary beam directions when describing the following two physicals effects: (i) the mean radiation pressure force stemming from a beam, hereafter referred for brevity as the radiation pressure force, (ii) the diffusion resulting from its fluctuations. We discuss these effects in case of a 2D optical molasses (2D OM), a Zeeman slower (ZS), a 2D magneto-optical trap (2D MOT), and a 3D MOT. The corresponding numerical setup is seen in Fig. 1(b), for reference.
@@ A. Radiation pressure force
$\textbf{F}_{tot}(\textbf{r},\textbf{v})=\textbf{F}_{\textsf{3M}}(\textbf{r},\textbf{v})+\textbf{F}_{\textsf{2M}}(\textbf{r},\textbf{v})+\textbf{F}_{\textsf{2O}}(\textbf{r},\textbf{v})+\textbf{F}_{\textsf{ZS}}(\textbf{r},\textbf{v})$ (A1)
@@
[…]
The unit vector ±$\pm\hat{\boldsymbol{\alpha}}^{\prime}_{\mathbbm{1}}$ denotes an arbitrary direction (for 2D MOT and 2D OM only two of the three subequations below are applicable, and for ZS only one):
±$\pm\hat{\boldsymbol{\mathsf{x}}}^{\prime}_{\mathbbm{1}}&=\pm(\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}),\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}),-\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}))\\ \hskip 3.0pt\pm\hat{\boldsymbol{\mathsf{y}}}^{\prime}_{\mathbbm{1}}&=\pm(-\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}),\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}),\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}))\\ \hskip 3.0pt\pm\hat{\boldsymbol{\mathsf{z}}}^{\prime}_{\mathbbm{1}}&=\pm(\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}}),-\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}}),\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}}))\end{aligned}$ (A4)
@@
The coefficient $p_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,Q}$ denotes the fraction of the ±$\pm{\boldsymbol{\alpha}}^{\prime}_{\mathbbm{1}}$ directed light of $\sigma^{-}$, $\pi$, or $\sigma^{+}$ polarization (described by $Q$) that drives the corresponding transition (described by $q$):
@@
$p_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,\sigma^{+}}(\textbf{r})$ =$=\begin{cases}\left(\frac{1}{2}\left[1+\frac{\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{B}(\textbf{r})}{|\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}||\textbf{B}(\textbf{r})|}\right]\right)^{2}\quad\quad\;\;,\quad q=\sigma^{+}\\ \left(\frac{1}{2}\left[1-\frac{\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{B}(\textbf{r})}{|\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}||\textbf{B}(\textbf{r})|}\right]\right)^{2}\quad\quad\;\;,\quad q=\sigma^{-}\\ 1-(p_{\alpha^{\prime},\sigma^{+};\mathbbm{1}}^{\pm,\sigma+}+p_{\alpha^{\prime},\sigma^{-};\mathbbm{1}}^{\pm,\sigma+})\;\;\;\,\quad\;\,,\quad q=\pi\end{cases}$ (A5)
$p_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,\sigma^{-}}(\textbf{r})$ =$=\begin{cases}\left(\frac{1}{2}\left[1-\frac{\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{B}(\textbf{r})}{|\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}||\textbf{B}(\textbf{r})|}\right]\right)^{2}\quad\quad\;\;,\quad q=\sigma^{+}\\ \left(\frac{1}{2}\left[1+\frac{\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{B}(\textbf{r})}{|\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}||\textbf{B}(\textbf{r})|}\right]\right)^{2}\quad\quad\;\;,\quad q=\sigma^{-}\\ 1-(p_{\alpha^{\prime},\sigma^{+};\mathbbm{1}}^{\pm,\sigma-}+p_{\alpha^{\prime},\sigma^{-};\mathbbm{1}}^{\pm,\sigma-})\;\;\;\;\quad\,\,,\quad q=\pi\end{cases}$
$p_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,\pi}(\textbf{r})$ =$=\begin{cases}\frac{1}{2}\left(1-p_{\alpha^{\prime},\pi;\mathbbm{1}}^{\pm,\pi}\right)\quad\quad\quad\quad\quad\quad\;,\quad q=\sigma^{+}\\ \frac{1}{2}\left(1-p_{\alpha^{\prime},\pi;\mathbbm{1}}^{\pm,\pi}\right)\quad\quad\quad\quad\quad\quad\;,\quad q=\sigma^{-}\\ \left(\frac{\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{B}(\textbf{r})}{|\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}||\textbf{B}(\textbf{r})|}\right)^{2}\quad\quad\quad\quad\quad\quad\,,\quad q=\pi\end{cases}$
@@
where $\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}=\pm k_{L}\hat{\boldsymbol{\alpha}}^{\prime}_{\mathbbm{1}}$ is the wavevector with the wavenumber $k_{L}=\omega_{L}/c$; and B is the total magnetic field experienced by the atom, written as the sum of the magnetic fields stemming from the different parts of the system. Their respective components are written below. (i) For the 3D MOT:
$B_{\mathsf{x_{\textsf{3M}}}}$ $=-\frac{B^{\prime}R}{2}\frac{(1+\frac{\eta^{2}}{4})^{5/2}}{3\eta}\frac{x_{\textsf{3M}}}{r_{\textsf{3M}}}$ (A6)
@@
×$\times\left(\frac{1}{\left(1+[\frac{r_{\textsf{3M}}}{R}-\frac{\eta}{2}]^{2}\right)^{3/2}}-\frac{1}{\left(1+[\frac{r_{\textsf{3M}}}{R}+\frac{\eta}{2}]^{2}\right)^{3/2}}\right)$
$B_{\mathsf{y_{\textsf{3M}}}}$ $=-\frac{B^{\prime}R}{2}\frac{(1+\frac{\eta^{2}}{4})^{5/2}}{3\eta}\frac{y_{\textsf{3M}}}{r_{\textsf{3M}}}$
×$\times\left(\frac{1}{\left(1+[\frac{r_{\textsf{3M}}}{R}-\frac{\eta}{2}]^{2}\right)^{3/2}}-\frac{1}{\left(1+[\frac{r_{\textsf{3M}}}{R}+\frac{\eta}{2}]^{2}\right)^{3/2}}\right)$
$B_{\mathsf{z_{\textsf{3M}}}}$ $=B^{\prime}R\frac{(1+\frac{\eta^{2}}{4})^{5/2}}{3\eta}$
×$\times\left(\frac{1}{\left(1+[\frac{z_{\textsf{3M}}}{R}-\frac{\eta}{2}]^{2}\right)^{3/2}}-\frac{1}{\left(1+[\frac{z_{\textsf{3M}}}{R}+\frac{\eta}{2}]^{2}\right)^{3/2}}\right)$
where $r_{\mathsf{3M}}=\sqrt{x_{\mathsf{3M}}^{2}+y_{\mathsf{3M}}^{2}+z_{\mathsf{3M}}^{2}}$ is the radial length with $x_{\mathsf{3M}},y_{\mathsf{3M}},z_{\mathsf{3M}}$ being the 3D MOT coordinates (respectively equal to $x,y,z$ minus the corresponding 3D MOT center coordinate); $B^{\prime}$ is the field gradient; $R$ is the coil radius; and $\eta$ is the ratio of the coil separation to the coil
@@
radius. It is well-known that the general anti-Helmholtz field involves elliptic integrals, and here a less complicated expression of its radial and axial components is provided. Note that the inclusion of $z_{\mathsf{3M}}$ in the radial components $B_{\mathsf{x_{\textsf{3M}}}}$ and $B_{\mathsf{y_{\textsf{3M}}}}$ results in them vanishing for distances far from the trap center. Also note that that the field approaches $B^{\prime}(-\frac{x_{\textsf{3M}}}{2},-\frac{y_{\textsf{3M}}}{2},z_{\textsf{3M}})$ when working on-axis and close to the trap center, usually considered in MOT descriptions.
@@
$B_{\mathsf{x_{\textsf{2M}}}}&=B^{\wedge,+}\times(3y_{\textsf{2M}}^{\wedge,+}x_{\textsf{2M}}^{\wedge,+})\\ &+B^{\wedge,-}\times(-3y_{\textsf{2M}}^{\wedge,-}x_{\textsf{2M}}^{\wedge,-})\\ &+B^{\vee,+}\times(3y_{\textsf{2M}}^{\vee,+}x_{\textsf{2M}}^{\vee,+})\\ &+B^{\vee,-}\times(-3y_{\textsf{2M}}^{\vee,-}x_{\textsf{2M}}^{\vee,-})\\ B_{\mathsf{y_{\textsf{2M}}}}&=B^{\wedge,+}\times(2[y_{\textsf{2M}}^{\wedge,+}]^{2}-[x_{\textsf{2M}}^{\wedge,+}]^{2}-[z_{\textsf{2M}}^{\wedge,+}]^{2})\\ &+B^{\wedge,-}\times(-2[y_{\textsf{2M}}^{\wedge,-}]^{2}+[x_{\textsf{2M}}^{\wedge,-}]^{2}+[z_{\textsf{2M}}^{\wedge,-}]^{2})\\ &+B^{\vee,+}\times(2[y_{\textsf{2M}}^{\vee,+}]^{2}-[x_{\textsf{2M}}^{\vee,+}]^{2}-[z_{\textsf{2M}}^{\vee,+}]^{2})\\ &+B^{\vee,-}\times(-2[y_{\textsf{2M}}^{\vee,-}]^{2}+[x_{\textsf{2M}}^{\vee,-}]^{2}+[z_{\textsf{2M}}^{\vee,-}]^{2})\\ B_{\mathsf{z_{\textsf{2M}}}}&=B^{\wedge,+}\times(3y_{\textsf{2M}}^{\wedge,+}z_{\textsf{2M}}^{\wedge,+})\\ &+B^{\wedge,-}\times(-3y_{\textsf{2M}}^{\wedge,-}z_{\textsf{2M}}^{\wedge,-})\\ &+B^{\vee,+}\times(3y_{\textsf{2M}}^{\vee,+}z_{\textsf{2M}}^{\vee,+})\\ &+B^{\vee,-}\times(-3y_{\textsf{2M}}^{\vee,-}z_{\textsf{2M}}^{\vee,-})\end{aligned}$ (A7)
@@
where $\wedge$” and $\vee$” in the superscripts refer to the magnets respectively above and below the center xy-plane of the 2D MOT, while ”$+$” and ”--” refer to them being respectively positively and negatively displaced in the x direction with respect to the 2D MOT center. $B^{*}=\frac{B_{0}V}{4\pi([x_{\textsf{2M}}^{*}]^{2}+[y_{\textsf{2M}}^{*}]^{2}+[z_{\textsf{2M}}^{*}]^{2})^{5/2}}$ is the field magnitude of the corresponding magnet, with $B_{0}$ being the remanent magnetization and $V$ the magnet volume; and $x_{\textsf{2M}}^{*}=x_{\textsf{2M}}-x^{*}$, $y_{\textsf{2M}}^{*}=y_{\textsf{2M}}-y^{*}$, $z_{\textsf{2M}}^{*}=z_{\textsf{2M}}-z^{*}$, where $x_{\textsf{2M}},y_{\textsf{2M}},z_{\textsf{2M}}$ are the 2D MOT coordinates, and $x^{*},y^{*},z^{*}$ are the magnet coordinates (”*” refers to either ”,$\wedge$,+∧ , +”, ”$\wedge$,-∧ , -”, ”,$\vee$,+∨ , +”, or ”$\vee$,-∨ , -”). We note that Eq. A7 has been obtained using the field equation for a magnetic dipole.
@@
(iii) For the 2D OM, the magnetic field is not present, and thus $\textbf{B}=\textbf{0}=(0,0,0)$. This will result in only the $\sigma^{+}$ and $\sigma^{-}$ transitions being driven according to Eq. A5.
(iv) For the ZS, we use a heuristic model based on segmen-
tation with linear functions:
$B_{\mathsf{x_{\textsf{ZS}}}}$ $=0$ (A8)
@@
$B_{\mathsf{y_{\textsf{ZS}}}}$ $=\sum_{n=1,2,...,n_{S}}\left[H(x_{\textsf{ZS}}-(n-1)L/n_{S})-H(x_{\textsf{ZS}}-nL/n_{S})\right]$
×[$\times\left[\left\{(B_{n+1}-B_{n})/(L/n_{S})\right\}\times\left\{(x_{\textsf{ZS}}-(n-1)L/n_{S}\right\}+B_{n})\right]$
$B_{\mathsf{z_{\textsf{ZS}}}}$ $=0$
@@
where the Heaviside functions $H$ provide a domain of truncation, with the outside field being zero; $x_{\textsf{ZS}}$ is the $\mathsf{x}$ coordinate of the ZS (obtained by subtracting $x$ with the corresponding ZS entrance coordinate); $B_{1,2,...,n_{S}+1}$ is the size of the magnetic field at a given location, with $B_{1}$ being the value at the ZS entrance and so forth until $B_{n_{S}+1}$, where $n_{S}$ is the amount of segments. We note that the magnetic field here is perpendicular to the non-rotated propagation direction of the Zeeman beam ($-\hat{\boldsymbol{\mathsf{x}}}$). For a two-level model, this would result in a vanishing radiation pressure force; on the other hand, as our model is sensitive to light polarization, the $\sigma^{+}$ and $\sigma^{-}$ transitions will be driven according to Eq. A5.

### v2

@@ V. Acknowledgments
We are grateful for contributions to the experimental setup by A. Rybicki and J. McDowell, and our cesium team for sharing their data [67]. This work was supported by an award from the W. M. Keck Foundation.
Appendix A
@@
We base our model on the hyperfine transition $F=0\rightarrow F^{\prime}=1$, faithful to the blue (or red) trapping descriptions of 88Sr. Here, each of the three Zeeman transitions $m_{0}=0\rightarrow m_{1}=-1,0,+1$ between the hyperfine levels is treated as an independent two-level system and driven by, respectively, $\sigma^{-}$, $\pi$, $\sigma^{+}$ polarized light (refer to Fig. 1(b) in Ref. [55]). This approximation is expected to hold only in the regime of low saturation, $s=\frac{I_{L}/I_{sat}}{1+\frac{4\Delta^{2}}{\Gamma^{2}}}\ll 1$, where $I_{L}$ is the laser intensity, $I_{sat}$ is the saturation intensity, $\Delta=\omega_{L}-\omega_{0}$ is the (angular) detuning of the laser frequency $\omega_{L}$ from the atomic transition $m_{0}=0\rightarrow m_{1}=0$ frequency $\omega_{0}$, and $\Gamma$ is the natural linewidth of the hyperfine transition.
We take into account arbitrary beam directions when describing the following two physicals effects: (i) the mean radiation pressure force stemming from a beam, hereafter referred to, for brevity, as the radiation pressure force, and (ii) the diffusion resulting from its fluctuations. We discuss these effects for the case of a 2D optical molasses (2D OM), a Zeeman slower (ZS), a 2D magneto-optical trap (2D MOT), and a 3D MOT. The corresponding numerical setup is seen in Fig. 2(b), for reference.
@@ A. Radiation pressure force
$\hskip-89.5pt\textbf{F}_{tot}(\textbf{r},\textbf{v})=\textbf{F}_{\textsf{3M}}(\textbf{r},\textbf{v})+\textbf{F}_{\textsf{2M}}(\textbf{r},\textbf{v})+\textbf{F}_{\textsf{2O}}(\textbf{r},\textbf{v})+\textbf{F}_{\textsf{ZS}}(\textbf{r},\textbf{v})$ (A1)
@@
In this equation, $\mathbbm{1}=\textsf{3M},\textsf{2M},\textsf{2O}$, or ZS (respectively referring to the 3D MOT, 2D MOT, 2D OM, or ZS); $c$ is the vacuum light speed; $\textbf{r}=(x,y,z)$ is the atom position; $\textbf{v}=(v_{\mathsf{x}},v_{\mathsf{y}},v_{\mathsf{z}})$ is the atom velocity; $q$ refers to the respective $\sigma^{-}$, $\pi$, $\sigma^{+}$ atomic transitions, with $Q$ referring to the corresponding beam polarizations; and the remaining quantities are defined as follows.
The unit vector $\pm\hat{\boldsymbol{\alpha}}^{\prime}_{\mathbbm{1}}$ denotes an arbitrary direction (for 2D MOT and 2D OM only two of the three subequations below are applicable, and for ZS only one):
$\hskip 3.0pt\begin{aligned} \pm\hat{\boldsymbol{\mathsf{x}}}^{\prime}_{\mathbbm{1}}&=\pm(\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}),\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}),-\text{sin}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}})\text{cos}(\phi_{\mathsf{x}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}))\\ \hskip 3.0pt\pm\hat{\boldsymbol{\mathsf{y}}}^{\prime}_{\mathbbm{1}}&=\pm(-\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}),\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}),\text{sin}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{cos}(\phi_{\mathsf{y}^{\prime};\mathbbm{1}}^{\pm,\mathsf{z}}))\\ \hskip 3.0pt\pm\hat{\boldsymbol{\mathsf{z}}}^{\prime}_{\mathbbm{1}}&=\pm(\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}}),-\text{sin}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}}),\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{x}})\text{cos}(\phi_{\mathsf{z}^{\prime};\mathbbm{1}}^{\pm,\mathsf{y}}))\end{aligned}$ (A4)
@@
The coefficient $p_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,Q}$ denotes the fraction of the $\pm{\boldsymbol{\alpha}}^{\prime}_{\mathbbm{1}}$ directed light of $\sigma^{-}$, $\pi$, or $\sigma^{+}$ polarization (described by $Q$) that drives the corresponding transition (described by $q$):
@@
$\hskip 11.0pt\begin{aligned} &p_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,\sigma^{+}}(\textbf{r})&&=\begin{cases}\left(\frac{1}{2}\left[1+\frac{\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{B}(\textbf{r})}{|\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}||\textbf{B}(\textbf{r})|}\right]\right)^{2}\quad\quad\;\;,\quad q=\sigma^{+}\\ \left(\frac{1}{2}\left[1-\frac{\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{B}(\textbf{r})}{|\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}||\textbf{B}(\textbf{r})|}\right]\right)^{2}\quad\quad\;\;,\quad q=\sigma^{-}\\ 1-(p_{\alpha^{\prime},\sigma^{+};\mathbbm{1}}^{\pm,\sigma+}+p_{\alpha^{\prime},\sigma^{-};\mathbbm{1}}^{\pm,\sigma+})\;\;\;\,\quad\;\,,\quad q=\pi\end{cases}\\ &p_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,\sigma^{-}}(\textbf{r})&&=\begin{cases}\left(\frac{1}{2}\left[1-\frac{\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{B}(\textbf{r})}{|\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}||\textbf{B}(\textbf{r})|}\right]\right)^{2}\quad\quad\;\;,\quad q=\sigma^{+}\\ \left(\frac{1}{2}\left[1+\frac{\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{B}(\textbf{r})}{|\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}||\textbf{B}(\textbf{r})|}\right]\right)^{2}\quad\quad\;\;,\quad q=\sigma^{-}\\ 1-(p_{\alpha^{\prime},\sigma^{+};\mathbbm{1}}^{\pm,\sigma-}+p_{\alpha^{\prime},\sigma^{-};\mathbbm{1}}^{\pm,\sigma-})\;\;\;\;\quad\,\,,\quad q=\pi\end{cases}\\ &p_{\alpha^{\prime},q;\mathbbm{1}}^{\pm,\pi}(\textbf{r})&&=\begin{cases}\frac{1}{2}\left(1-p_{\alpha^{\prime},\pi;\mathbbm{1}}^{\pm,\pi}\right)\quad\quad\quad\quad\quad\quad\;,\quad q=\sigma^{+}\\ \frac{1}{2}\left(1-p_{\alpha^{\prime},\pi;\mathbbm{1}}^{\pm,\pi}\right)\quad\quad\quad\quad\quad\quad\;,\quad q=\sigma^{-}\\ \left(\frac{\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}\cdot\textbf{B}(\textbf{r})}{|\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}||\textbf{B}(\textbf{r})|}\right)^{2}\quad\quad\quad\quad\quad\quad\,,\quad q=\pi\end{cases}\end{aligned}$ (A5)
@@
where $\textbf{k}_{\alpha^{\prime};\mathbbm{1}}^{\pm}=\pm k_{L}\hat{\boldsymbol{\alpha}}^{\prime}_{\mathbbm{1}}$ is the wavevector with the wavenumber $k_{L}=\omega_{L}/c$; and B is the total magnetic field experienced by the atom, written as the sum of the magnetic fields stemming from the different parts of the system. Their respective components are written below.
(i) For the 3D MOT:
@@
$\normalsize{}\small\hskip-100.0pt\begin{aligned} B_{\mathsf{x_{\textsf{3M}}}}&=\frac{B^{\prime}R}{2}f\frac{x_{\textsf{3M}}}{r_{\textsf{3M}}}\times\left(\frac{1}{\left(1+[\frac{r_{\textsf{3M}}}{R}+\frac{\eta}{2}]^{2}\right)^{3/2}}-\frac{1}{\left(1+[\frac{r_{\textsf{3M}}}{R}-\frac{\eta}{2}]^{2}\right)^{3/2}}\right)\\ B_{\mathsf{y_{\textsf{3M}}}}&=\frac{B^{\prime}R}{2}f\frac{y_{\textsf{3M}}}{r_{\textsf{3M}}}\times\left(\frac{1}{\left(1+[\frac{r_{\textsf{3M}}}{R}+\frac{\eta}{2}]^{2}\right)^{3/2}}-\frac{1}{\left(1+[\frac{r_{\textsf{3M}}}{R}-\frac{\eta}{2}]^{2}\right)^{3/2}}\right)\\ B_{\mathsf{z_{\textsf{3M}}}}&=-B^{\prime}Rf\times\left(\frac{1}{\left(1+[\frac{z_{\textsf{3M}}}{R}+\frac{\eta}{2}]^{2}\right)^{3/2}}-\frac{1}{\left(1+[\frac{z_{\textsf{3M}}}{R}-\frac{\eta}{2}]^{2}\right)^{3/2}}\right)\end{aligned}$ (A6)
@@
where $r_{\mathsf{3M}}=\sqrt{x_{\mathsf{3M}}^{2}+y_{\mathsf{3M}}^{2}+z_{\mathsf{3M}}^{2}}$ is the total radial distance with $x_{\mathsf{3M}},y_{\mathsf{3M}},z_{\mathsf{3M}}$ being the 3D MOT coordinates (respectively equal to $x,y,z$ minus the corresponding 3D MOT center coordinate); $B^{\prime}$ is the field gradient; $R$ is the coil radius; $\eta$ is the ratio of the coil separation to the coil radius; and the coefficient $f=\frac{\left(1+\frac{\eta^{2}}{4}\right)^{5/2}}{3\eta}$. It is well known that the general anti-Helmholtz field involves elliptic integrals, and here a less complicated form of its radial and axial components is assumed. Note that we take into account the radial components vanishing for distances far from the trap center. Also note that the field approaches $B^{\prime}\left(\frac{x_{\textsf{3M}}}{2},\frac{y_{\textsf{3M}}}{2},-z_{\textsf{3M}}\right)$ when working on-axis and close to the trap center, usually considered in MOT descriptions.
@@
$\hskip-115.0pt\begin{aligned} B_{\mathsf{x_{\textsf{2M}}}}&=B^{\wedge,+}\times(3y_{\textsf{2M}}^{\wedge,+}x_{\textsf{2M}}^{\wedge,+})\\ &+B^{\wedge,-}\times(-3y_{\textsf{2M}}^{\wedge,-}x_{\textsf{2M}}^{\wedge,-})\\ &+B^{\vee,+}\times(3y_{\textsf{2M}}^{\vee,+}x_{\textsf{2M}}^{\vee,+})\\ &+B^{\vee,-}\times(-3y_{\textsf{2M}}^{\vee,-}x_{\textsf{2M}}^{\vee,-})\\ B_{\mathsf{y_{\textsf{2M}}}}&=B^{\wedge,+}\times(2[y_{\textsf{2M}}^{\wedge,+}]^{2}-[x_{\textsf{2M}}^{\wedge,+}]^{2}-[z_{\textsf{2M}}^{\wedge,+}]^{2})\\ &+B^{\wedge,-}\times(-2[y_{\textsf{2M}}^{\wedge,-}]^{2}+[x_{\textsf{2M}}^{\wedge,-}]^{2}+[z_{\textsf{2M}}^{\wedge,-}]^{2})\\ &+B^{\vee,+}\times(2[y_{\textsf{2M}}^{\vee,+}]^{2}-[x_{\textsf{2M}}^{\vee,+}]^{2}-[z_{\textsf{2M}}^{\vee,+}]^{2})\\ &+B^{\vee,-}\times(-2[y_{\textsf{2M}}^{\vee,-}]^{2}+[x_{\textsf{2M}}^{\vee,-}]^{2}+[z_{\textsf{2M}}^{\vee,-}]^{2})\\ B_{\mathsf{z_{\textsf{2M}}}}&=B^{\wedge,+}\times(3y_{\textsf{2M}}^{\wedge,+}z_{\textsf{2M}}^{\wedge,+})\\ &+B^{\wedge,-}\times(-3y_{\textsf{2M}}^{\wedge,-}z_{\textsf{2M}}^{\wedge,-})\\ &+B^{\vee,+}\times(3y_{\textsf{2M}}^{\vee,+}z_{\textsf{2M}}^{\vee,+})\\ &+B^{\vee,-}\times(-3y_{\textsf{2M}}^{\vee,-}z_{\textsf{2M}}^{\vee,-})\end{aligned}$ (A7)
@@
where $\wedge$” and $\vee$” in the superscripts refer to the magnets respectively above and below the center xy plane of the 2D MOT, while “++” and “-” refer to them being respectively positively and negatively displaced in the x direction with respect to the 2D MOT center. $B^{*}=\frac{B_{0}V}{4\pi([x_{\textsf{2M}}^{*}]^{2}+[y_{\textsf{2M}}^{*}]^{2}+[z_{\textsf{2M}}^{*}]^{2})^{5/2}}$ is the field magnitude of the corresponding magnet, with $B_{0}$ being the remanent magnetization and $V$ the magnet volume; and $x_{\textsf{2M}}^{*}=x_{\textsf{2M}}-x^{*}$, $y_{\textsf{2M}}^{*}=y_{\textsf{2M}}-y^{*}$, $z_{\textsf{2M}}^{*}=z_{\textsf{2M}}-z^{*}$, where $x_{\textsf{2M}},y_{\textsf{2M}},z_{\textsf{2M}}$ are
the 2D MOT coordinates, and $x^{*},y^{*},z^{*}$ are the magnet coordinates (“*” refers to either $\wedge,+$”, $\wedge,$-”, $\vee,+$”, or $\vee,$-”). We note that Eq. A7 has been obtained using the field equation for a magnetic dipole.
@@
(iii) For the 2D OM, the magnetic field is not present [i.e., $\textbf{0}=(0,0,0)$]. This will result in only the $\sigma^{+}$ and $\sigma^{-}$ transitions being driven according to Eq. A5.
(iv) For the ZS, we use a heuristic model based on segmentation with linear functions:
@@
$\vskip-3.0pt\hskip 5.0pt\begin{aligned} B_{\mathsf{x_{\textsf{ZS}}}}&=0\\ B_{\mathsf{y_{\textsf{ZS}}}}&=\sum_{n=1,2,...,n_{S}}\left\{H(x_{\textsf{ZS}}-(n-1)L/n_{S})-H(x_{\textsf{ZS}}-nL/n_{S})\right\}\\ &\times\left\{\left[(B_{n+1}-B_{n})/(L/n_{S})\right]\times\left[x_{\textsf{ZS}}-(n-1)L/n_{S}\right]+B_{n}\right\}\\ B_{\mathsf{z_{\textsf{ZS}}}}&=0\end{aligned}$ (A8)
@@
where the Heaviside functions $H$ provide a domain of truncation, with the outside field being zero; $x_{\textsf{ZS}}$ is the $\mathsf{x}$ coordinate of the ZS (obtained by subtracting $x$ with the corresponding ZS entrance coordinate); $L$ is the length of the ZS; $B_{1,2,...,n_{S}+1}$ is the size of the magnetic field at a given location, with $B_{1}$ and $B_{n_{S}+1}$ being the values at the ZS entrance and exit, respectively, where $n_{S}$ is the amount of segments. We note that the magnetic field here is perpendicular to the non rotated propagation direction of the Zeeman beam ($\hat{\boldsymbol{\mathsf{x}}}$). For a two-level model, this would result in a vanishing radiation pressure force; on the other hand, as our model is sensitive to light polarization, the $\sigma^{+}$ and $\sigma^{-}$ transitions will be driven according to Eq. A5.